Logiweb bibliography of factorial

Up Help

[0] factorial

[1] base

The pyk compiler, version 0.grue.20050502 by Klaus Grue GRD-2005-06-29.UTC:01:03:50.943925 = MJD-53550.TAI:01:04:22.943925 = LGT-4626723862943925e-6